Provide the value of arccos(sin(14π/13)) without the use of a calculator


1
Expert's answer
2021-05-10T13:05:10-0400

using

"cos(\frac{\pi}{2}-\theta)=sin(\theta)"


"arc \space cos (sin(\frac{14\pi}{13}))"


"arc\space cos (sin(\frac{14\pi}{13}))=arc\space cos(cos(\frac{\pi}{2}-\frac{14\pi}{13}))"


"=arc\space cos(cos(\frac{13\pi-28\pi}{26}))"


"=arc\space cos (cos(\frac{-15\pi}{26}))"


"=arc\space cos(cos(\frac{15\pi}{26}))"


"=\frac{15\pi}{26}"


"=1.81245"